# GenCeption: Evaluate Multimodal LLMs with Unlabeled Unimodal Data

### v3

@@ 3 Experiments
Table 3: The impact of different image encoders, DALL·E 3 (Dalle3) vs. Imagen 2 (Imgn2), on GC@$1$ score. Best results per configuration and category (over different MLLMs) are bolded.
3.1 Quantitative results
We partition the 14 MME categories into two groups based on content type: visual-intensive (10 categories: existence, count, position, color, poster, celebrity, scene, landmark, artwork, and commonsense reasoning) and textual-intensive (4 categories: code reasoning, numerical calculation, text translation, and OCR). GC@$3$ scores on the MMECeption benchmark and accuracy on the original MME benchmark are reported per category and as aggregations in Table 2. Additionally, we include the scores and ranks of all evaluated MLLMs on the OpenCompass OpenCompass (2023), MME Fu et al. (2023), HallusionBench Liu et al. (2023a), MMStar Chen et al. (2024), SeedBench Li et al. (2023a), and AI2D Kembhavi et al. (2016) leaderboards. Notably, Gemini1.5-Pro leads our rankings, followed by GPT-4o, GPT-4v, Claude3-Opus, mPLUG-Owl2, and LLaVA-13B/7B. The GenCeption method shows robustness to the similarity metric used, as the overall ranking remains identical when using cosine similarity or FID distance for calculating GC@$T$ scores.
@@ 3.1 Quantitative results
Table 4: The performance of VLLMs and humans on the GC@$1$ metric, evaluated using 5 randonly drawn images per sample/image category. the best performance achieved by an MLLM is underlined.
@@
Figure 2 presents a correlation matrix among GC@$T$ scores and the benchmarks mentioned above, where the overall GC@$T$ scores are averaged over the GC@$T$ scores of all MME categories. The strong correlations with the OpenCompass scores incorporating the results of multiple meticulously crafted benchmarks indicate that MMECeption provides a comprehensive evaluation that may complement existing benchmarks. Further, GenCeption appears to effectively measure an MLLM’s tendency to hallucinate, as demonstrated by the strong correlations with HallusionBench. While these observations are further emphasized by the correlations with MMStar and AI2D, the only moderate correlations with MME and SEEDBench provide more nuanced insights. As MME displays these moderate correlations also with the other benchmarks, it can be reasoned that it measures dimensions supplement to those measured by other benchmarks and GenCeption. SEEDBench on the other hand correlates strongly with other benchmarks, but only moderately with GC@$T$ scores. This indicates that SEEDBench measures aspects that are also measured by other benchmarks, but fail to be captured by GenCeption. Future research could focus on identifying these aspects to potentially incorporate them into GenCeption.
@@
Table 3 compares GC@$1$ scores using different image generators, OpenAI’s DALL·E 3 Ramesh et al. (2021) and Google DeepMind’s Imagen2 DeepMind (2023). Independent of image generator used, the rankings remain unchanged, except that on visual-intensive samples only, Claude3-Opus scores equally with GPT-4V. This provides evidence that even though DALL·E 3, GPT-4o, and GPT-4V were developed and trained by OpenAI, neither of OpenAI’s models has an advantage over non-OpenAI MLLMs.
@@
Table 4 shows human performance on a subset of 5 randomly drawn images per category compared to the MLLM performance on the same subset of samples. It can be observed that humans outperform all MLLMs, with especially strong differences in performance for the text-intensive categories. The worst performance, relative to humans, is achieved on the code reasoning and text translation categories, the former containing images of code snippets and the latter of phrases written in simplified Chinese characters. The relatively best performance by MLLMs is achieved on the scene and artwork categories, which contain every-day life photos and popular artworks. This demonstrates that there is still substantial space for performance improvement, and that compared to humans, MLLMs still lack relevant competences. It must be noted that human performance here does not constitute an upper bound in possible scores to achieve, and that future generations of MLLMs may well outperform humans.
@@ 3.2 Qualitative Results
We qualitatively inspect our results by visualizing generated images together with their cosine similarity and GC@$T$ scores for two seed images across different categories, as shown in Figure 3. This visualization highlights the correlation between these scores and the visual characteristics of the images relative to the seed image. A key observation is that later iterations show an increased tendency to produce imagery deviating from the seed image, as indicated by lower GC@$T$ scores. This serves as an additional qualitative validation of the GenCeption method and the MMECeption benchmark, as using MLLMs scoring higher on the MMECeption benchmark results in generated images that preserve more information from the seed image. For a wider range of examples across MME image categories and corresponding descriptions from each evaluated VLLM, readers are referred to Appendix A.
@@ 4 Discussion and Future Directions
This study validates the GenCeption method focusing on the visual modality only. Future research is invited to adapt GenCeption to other modalities, such as audio, video, and graphs. For instance, the framework can be initiated with a dataset of audio samples, and MLLMs can iteratively generate and describe the audio content. Similarly, for video and graph data, the process can involve generating textual descriptions of short video clips or graph structures and their recreation. While the core iterative process of GenCeption remains applicable, these extensions require modality-specific generation and embedding models.
@@ 5 Conclusion
In this paper, we introduce GenCeption to enhance the evaluation of rapidly evolving Multimodal Language Models (MLLMs). The GenCeption method addresses key limitations of existing MLLM benchmarks, such as costly data annotation, quick benchmark saturation, training data contamination, leading questions, and the illusion of emergent abilities. Being adaptable to different modalities, the GenCeption method can deliver value as a unified MLLM evaluation method that complements existing MLLM benchmarks.
@@ Appendix A GenCeption Demonstration
Appendix B Limitations and Societal Impact
The limitations, outlined in Sections 3 and 4, primarily pertain to our initial experimental focus on image-based experiments, excluding other modalities. A critical assumption is the minimal influence of stochastic variability in image generation and MLLM text generation processes. While we have not delved into ethical risks, our framework’s purpose – to assess inter-modality semantic drift and susceptibility to hallucination in MLLMs—is clearly articulated. Societally, the exclusive use of the English language in GenCeption experiments may inadvertently marginalize non-English-speaking user groups.
Appendix C Dataset and Reproducibility
@@
In our study, we adopt several state-of-the-art models to facilitate our experiments, including Gemini1.5-Pro, GPT-4o, GPT-4V, Claude3-Opus, LLaVa-13B, LLaVa-7B, and mPLUG-Ow12 for text description generation, ViT for image embedding, and DALL·E 3 and Imagen2 for image generation, adhering to default parameter settings as outlined in their original specifications. As we only evaluated existing models but did not train new models, no hyperparamter tuning was applicaple. The text descriptions generated by GPT-4V are obtained through API calls, while experiments involving the other models are conducted on A100 GPUs, totaling approximately 96 GPU hours. Image generation was also performed via a call to OpenAI’s DALL-E 3 API. To compute the GC@$T$ metric, we employ the cosine similarity metric from the Scikit-learn library (Version 1.4.0).
Appendix D Human Annotators

### v4

@@ 3.1 Quantitative results
Table 4: The impact of different image encoders, DALL·E 3 (Dalle3) vs. Imagen 2 (Imgn2), on GC@$1$ score. Best results per configuration and category (over different VLLMs) are bolded.
@@
Table 5: The performance of VLLMs and humans on the GC@$1$ metric, evaluated using 5 randomly drawn images per sample/image category. The best performance achieved by an VLLM is underlined.
@@
Figure 2 presents a correlation matrix among GC@$T$ scores and the benchmarks mentioned above, where the overall GC@$T$ scores are averaged over the GC@$T$ scores of all MME categories. The strong correlations with the OpenCompass scores incorporating the results of multiple meticulously crafted benchmarks indicate that MMECeption provides a comprehensive evaluation that may complement existing benchmarks. Further, GenCeption appears to effectively measure a VLLM’s tendency to hallucinate, as demonstrated by the strong correlations with HallusionBench. While these observations are further emphasized by the correlations with MMStar and AI2D, the only moderate correlations with MME and SEEDBench provide more nuanced insights. As MME displays these moderate correlations also with the other benchmarks, it can be reasoned that it measures dimensions supplement to those measured by other benchmarks and GenCeption. SEEDBench on the other hand correlates strongly with other benchmarks, but only moderately with GC@$T$ scores. This indicates that SEEDBench measures aspects that are also measured by other benchmarks, but fail to be captured by GenCeption. Future research could focus on identifying these aspects to potentially incorporate them into GenCeption.
One of the key strengths of GenCeption lies in its annotation-free evaluation methodology, a concept also reflected in emerging evaluation methods such as CrossCheckGPT  Sun et al. (2024). CrossCheckGPT ranks hallucinations by evaluating the consistency of outputs across independent MLLM models. In 3, we analyze the correlation of CrossCheckGPT with GenCeption, MME, and HallusionBench scores. The results show strong correlations between CrossCheckGPT and both GenCeption and HallusionBench, affirming its capability to capture key evaluative metrics. Notably, CrossCheckGPT exhibits a weaker correlation with MME, which is likely because the GenCeption benchmark is developed using MME image samples, making it inherently more aligned with the MME framework.
GC@$T$ scores, as defined in Equation (3), are weighted by a temporal factor $t$. To examine the impact of this weighting, we conducted an ablation study where the weighting mechanism was removed, effectively transforming GC@$T$ into $s^{(T)}$. Table 3 demonstrates that $s^{(3)}$ retains a high correlation with GC@3, yet its correlation with MME diminishes compared to the weighted version, while its alignment with HallusionBench remains consistent. Furthermore, unweighted scores correlate with MME in a uniform manner across different iterations, whereas the weighted scores show a progressive increase in correlation with MME as more iterations are applied. This indicates that temporal weighting amplifies later iterations’ influence, emphasizing cumulative semantic shifts captured by MME’s iterative design. Generally, stronger correlation with MME is desirable as it validates the alignment between GenCeption’s metrics and an established benchmark, reinforcing GenCeption’s ability to assess iterative semantic coherence effectively and reliably.
@@
Table 4 compares GC@$1$ scores using different image generators, OpenAI’s DALL·E 3 (Ramesh et al., 2021) and Google DeepMind’s Imagen2 (DeepMind, 2023). Independent of image generator used, the rankings remain unchanged, except that on visual-intensive samples only, Claude3-Opus scores equally with GPT-4V. This provides evidence that even though DALL·E 3, GPT-4o, and GPT-4V were developed and trained by OpenAI, neither of OpenAI’s models has an advantage over non-OpenAI VLLMs.
@@
Table 5 shows human performance on a subset of 5 randomly drawn images per category compared to the VLLM performance on the same subset of samples. It can be observed that humans outperform all VLLMs, with especially strong differences in performance for the text-intensive categories. The worst performance, relative to humans, is achieved on the code reasoning and text translation categories, the former containing images of code snippets and the latter of phrases written in simplified Chinese characters. The relatively best performance by VLLMs is achieved on the scene and artwork categories, which contain every-day life photos and popular artworks. This demonstrates that there is still substantial space for performance improvement, and that compared to humans, VLLMs still lack relevant competences. It must be noted that human performance here does not constitute an upper bound in possible scores to achieve, and that future generations of VLLMs may well outperform humans.
@@ 3.2 Qualitative Results
We qualitatively inspect our results by visualizing generated images together with their cosine similarity and GC@$T$ scores for two seed images across different categories, as shown in Figure 3. This visualization highlights the correlation between these scores and the visual characteristics of the images relative to the seed image. A key observation is that later iterations show an increased tendency to produce imagery deviating from the seed image, as indicated by lower GC@$T$ scores. This serves as an additional qualitative validation of the GenCeption method and the MMECeption benchmark, as using VLLMs scoring higher on the MMECeption benchmark results in generated images that preserve more information from the seed image. For a wider range of examples across MME image categories and corresponding descriptions from each evaluated VLLM, readers are referred to A.
@@ 4 Discussion and Future Directions
This study validates the GenCeption method with a focus on the visual modality primarily because (1) VLLMs are the most widely used and readily available MLLMs on the market, and (2) image generation and embedding tools have reached a mature and highly commercialized stage compared to other modalities. However, GenCeption is designed to be modality-agnostic. The same iterative procedure (i.e., describing a unimodal sample and then re-generating it from the description) can, in principle, be applied to other non-text modalities like audio and video. The requirement is that (1) a generation model exists for the given modality, and (2) there is a suitable encoder to quantify the similarity between the original sample and the regenerated one. Moreover, recent advancements have introduced multimodal LLMs capable of both generating and interpreting multiple modalities simultaneously, such as Show-o (Xie et al., 2024), Emu3 (Wang et al., 2024), and JanusPro (Chen et al., 2025). In these cases, GenCeption could leverage the same MLLM for both description and generation tasks, serving as a particularly valuable approach for directly measuring modality consistency within such unified multimodal systems.
Future research is invited to adapt GenCeption to other non-text modalities, such as audio, video, and graphs. For instance, the framework can be initiated with a dataset of audio samples, and MLLMs can iteratively generate and describe the audio content. Similarly, for video and graph data, the process can involve generating textual descriptions of short video clips or graph structures and their recreation. While the core iterative process of GenCeption remains applicable, these extensions require careful exploration of modality-specific generation and embedding models.
@@ 5 Conclusion
In this paper, we introduce GenCeption to enhance the evaluation of rapidly evolving Multimodal Language Models (MLLMs). The GenCeption method attempts to address key limitations of existing MLLM benchmarks, such as costly data annotation, leading questions, the illusion of emergent abilities, and, as it allows to use newly created images without annotation, training data contamination. Further, it is expected to result in slower benchmark saturation. Being adaptable to different modalities, the GenCeption method can deliver value as a unified MLLM evaluation method that complements existing MLLM benchmarks.
@@ Appendix A GenCeption Demonstration
Figure 4: Example seed images from the visually (Figure 5) and textually (Figure 6) intensive groups, along with their associated metadata.
Appendix B Dataset and Reproducibility
@@
Figure 5: Illustration of a 3-iteration GenCeption procedure run on a visual-intensive image (from “existence” category) to evaluate 7 VLLMs. Each iteration $t$ shows the generated image $\mathbf{X}^{(t)}$, the description $\mathbf{Q}^{(t)}$ of the preceding image $\mathbf{X}^{(t-1)}$, and the similarity score $s^{(t)}$ relative to $\mathbf{X}^{(0)}$. The GC@$3$ metric for each VLLM is also presented. Hallucinated elements within descriptions $\mathbf{Q}^{(1)}$ and $\mathbf{Q}^{(2)}$ as compared to the seed image are indicated with red underlining.
Figure 6: Illustration of a 3-iteration GenCeption procedure run on a textual-intensive image (from “code reasoning” category) to evaluate 7 VLLMs. Each iteration $t$ shows the generated image $\mathbf{X}^{(t)}$, the description $\mathbf{Q}^{(t)}$ of the preceding image $\mathbf{X}^{(t-1)}$, and the similarity score $s^{(t)}$ relative to $\mathbf{X}^{(0)}$. The GC@$3$ metric for each VLLM is also presented. Hallucinated elements within descriptions $\mathbf{Q}^{(1)}$ as compared to the seed image are indicated with red underlining.
In our study, we adopt several state-of-the-art models to facilitate our experiments, including Gemini1.5-Pro, GPT-4o, GPT-4V, Claude3-Opus, LLaVa-13B, LLaVa-7B, and mPLUG-Ow12 for text description generation, ViT for image embedding, and DALL·E 3 and Imagen2 for image generation, adhering to default parameter settings as outlined in their original specifications. As we only evaluated existing models but did not train new models, no hyperparamter tuning was applicaple. The text descriptions generated by GPT-4V/4o, Claude3 and Gemini1.5 are obtained through API calls, while experiments involving the other models are conducted on A100 GPUs, totaling approximately 96 GPU hours. Image generation was also performed via a call to OpenAI’s DALL-E 3 API. To compute the GC@$T$ metric, we employ the cosine similarity metric from the Scikit-learn library (Version 1.4.0).
Appendix C Human Annotators